Loss amount, X, follows a gamma distribution. You are given E[X] = 8, Var[X] = 32. Calculate the probability that the loss amount exceeds 3, Pr[X>3].
1
Expert's answer
2020-12-08T07:48:57-0500

"\text{For gamma distribution}, \\
E(x)=\frac{n}{\lambda}, V(x)=\frac{n}{\lambda^2},then,\\
\frac{n}{\lambda}=8........(i)\\
\frac{n}{\lambda^2}=32.....(ii)\\
\text{By dividing\;(i)by(ii)}\\
\therefore \lambda=0.25, n=2\\
f(x)=0.0625xe^{-0.25x}, x\geq 0\\
Pr[X>3]=\int^{\infty}_{3}0.0625xe^{-0.25x}dx\\
=0.0625[\frac{-xe^{-0.25x}}{0.25}-\frac{e^{-0.25x}}{0.0625}]^\infty_3\\
=0.0625[0-(\frac{-3e^{-0.25(3)}}{0.25}-\frac{e^{-0.25(3)}}{0.0625})]\\
\approx 0.827"